### v1

# A note on Hilbert transform over lattices of $\mathrm{PSL}_{2}(\mathbb{C})$

Abstract.
In a recent work by González-Pérez, Parcet and Xia, the boundedness over non-commutative $L_{p}$-spaces of an analogue of the Hilbert transform was studied. This analogue is defined as a Fourier multiplier with symbol $m\colon\mathrm{PSL}_{2}(\mathbb{C})\to\mathbb{R}$ arising from the action by isometries of $\mathrm{PSL}_{2}(\mathbb{C})$ on the $3$-dimensional hyperbolic space $\mathbb{H}^{3}$. More concretely, $m$ is a lifting of the function on $\mathbb{H}^{3}$ that takes values ±$\pm 1$ in the two regions formed by dividing the space by a hyperbolic plane. The boundedness of $T_{m}$ on $L_{p}(\mathcal{L}\mathrm{PSL}_{2}(\mathbb{C}))$ for $p\neq 2$ was disproved by Parcet, de la Salle and Tablate. Nevertheless, we will show that this Fourier multiplier is bounded when restricted to the arithmetic lattices $\mathrm{PSL}_{2}(\mathbb{Z}[\sqrt{-n}])$, solving a question left open by González-Pérez, Parcet and Xia.
The author was partially supported by pre-doctoral scholarship PRE2020-093245, the Severo Ochoa Grant CEX2023-001347-5(MICIU), PIE2023-50E106(CSIC), and PID2022-141354NB-I00(MICINN)
Introduction
The boundedness problem for Fourier multipliers on $L_{p}$-spaces has always played a central role in harmonic analysis. One of the most studied examples is the Hilbert transform, defined as $\widehat{Hf}(\xi)=i\,\mathrm{sign}(\xi)\widehat{f}(\xi)$ for $f\in L_{2}(\mathbb{R})$. Although $H$ was already known to be bounded in $L_{p}(\mathbb{R})$ for $1<p<\infty$, in 1955 Cotlar [Cot55] gave a very simple proof of this fact using the following identity:
$|Hf|^{2}=2H(f\cdot Hf)-H(H(|f|^{2})).$ (Classical Cotlar)
This is known nowadays as the Cotlar identity. His proof uses that $H$ is bounded in $L_{2}(\mathbb{R})$ and that, by a recursive use of (Classical Cotlar), it also must be bounded in every $p=2^{k}$ for $k\geq 1$. Interpolation and the fact that $H$ is self-adjoint complete the proof.
Mei and Ricard [MR17] introduced the Cotlar identity in the non-commutative setting in order to study Hilbert transforms over free groups and amalgamated free products of von Neumann algebras. In the recent work of González-Pérez, Parcet and Xia [GPPX22] the authors developed a systematic approach to study Cotlar identities for Fourier multipliers in non-Abelian groups. Let $G$ be an unimodular group, $\mathcal{L}G$ the von Neumann algebra of $G$ and $G_{0}\subset G$ an open subgroup. Consider $m\colon G\to\mathbb{C}$ a symbol on $G$ and $T_{m}$ the corresponding Fourier multiplier on $\mathcal{L}G$. Then the formula:
$(m(g^{-1})-m(h))(m(gh)-m(g))=0,\quad\text{for all }g\in G\setminus G_{0},\,h\in G,$ (Cotlar)
is a translation of (Classical Cotlar) for $T_{m}$ in terms of its symbol. The main result in [GPPX22] states that any $m$ which is bounded, left $G_{0}$-invariant and verifies (Cotlar) defines a bounded multiplier in $L_{p}(\mathcal{L}\,G)$ for all $1<p<\infty$.
The subgroup $G_{0}$ represents a set in which the Cotlar identity may fail. In the argument, this failure is balanced by the invariance of $m$ respect $G_{0}$. Therefore this formulation of the theorem allows more flexibility in terms of the multiplier than the original one. However, the hypothesis of invariance can be relaxed even further. If $\chi\colon G_{0}\to\mathbb{T}^{1}$ is a character, it is enough for the result to hold that $m$ verifies:
$m(gh)=\chi(g)m(h)\quad\text{for all }g\in G_{0},\,h\in G.$
We say in this case that $m$ is left $(G_{0},\chi)$-invariant, and of course the $G_{0}$-invariance is recovered when $\chi$ is the trivial character.
Hilbert transform in $\mathrm{PSL}_{2}(\mathbb{C})$
Recall that $\mathrm{PSL}_{2}(\mathbb{C})$, which is the quotient of the $2\times 2$ complex matrices with determinant $1$ by its center, can be identified with the isometry group of the three dimensional hyperbolic space $\mathbb{H}^{3}$. This identification can be make explicit in various ways. Here we give one using the upper-space model of $\mathbb{H}^{3}$ and quaternions. Let $i,j,k$ denote the usual three quaternionic units, and let’s define:
$\mathbb{H}^{3}=\{a+bi+cj\colon a,b,c\in\mathbb{R},\,c>0\}.$
Doing so, $\mathbb{H}^{3}$ is exactly the subspace $\mathbb{C}+\mathbb{R}_{>0}j$ of the quaternions. Now, for a given $\omega\in\mathbb{H}^{3}$ we set:
$g\cdot\omega=(a\omega+b)(c\omega+d)^{-1},\quad\text{for }g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in\mathrm{PSL}_{2}(\mathbb{C}).$
This is a well-defined action of $\mathrm{PSL}_{2}(\mathbb{C})$ on $\mathbb{H}^{3}$, and it is an action by isometries when equipping $\mathbb{H}^{3}$ with the usual Riemannian structure.
On the other hand, a group $G$ acting on a set $X$ induces a multiplier on $G$ as follows: first choose a point $x_{0}\in X$ and two disjoint subsets $X^{+},X^{-}\subset X$. Let $m$ be the map $m\colon G\to\mathbb{C}$ defined for each $g\in G$ as:
$m(g)=\begin{cases}1&\text{if }g\cdot x_{0}\in X^{+},\\ -1&\text{if }g\cdot x_{0}\in X^{-},\\ 0&\text{otherwise}.\end{cases}$
Even if the final multiplier depends on $x_{0}$ and also on the partition given by $X^{+}$ and $X^{-}$, the boundedness of the multiplier is preserved by changing $x_{0}$ for any other point in the same $G$-orbit or using the sets $\{g\cdot X^{+},g\cdot X^{-}\}$, with $g\in G$, instead of $\{X^{+},X^{-}\}$. Back to the action of $\mathrm{PSL}_{2}(\mathbb{C})$ on the hyperbolic space, we are choosing the base point in $\mathbb{H}^{3}$ given by $j$ in our quaternionic parametrization, and the following partition:
+$\mathbb{H}^{3}_{+}=\{\omega\in\mathbb{H}^{3}\colon\operatorname{Re}(\omega)>0\},\quad\text{and}\quad\mathbb{H}^{3}_{-}=\{\omega\in\mathbb{H}^{3}\colon\operatorname{Re}(\omega)<0\}.$
This procedure induces a symbol $m$ in $\mathrm{PSL}_{2}(\mathbb{C})$ that is explicitly given by:
$m(g)=\mathrm{sign}(\operatorname{Re}(a\overline{c}+b\overline{d})),\quad\text{with }g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in\mathrm{PSL}_{2}(\mathbb{C}).$ (1)
The dividing frontier $\mathbb{H}^{3}\smallsetminus(\mathbb{H}^{3}_{+}\cup\mathbb{H}^{3}_{-})=\{\omega\in\mathbb{H}^{3}\colon\operatorname{Re}(\omega)=0\}$ is a hyperbolic plane, which determines the symbol $m$ up to a sign. Since the action of $\mathrm{PSL}_{2}(\mathbb{C})$ is transitive both on points and hyperbolic planes in $\mathbb{H}^{3}$, the boundedness of the multiplier defined by $m$ on $L_{p}(\mathcal{L}\,\mathrm{PSL}_{2}(\mathbb{C}))$ will remain the same under any other choice of the kind. Also, it is worth noticing that $m$ is easily shown invariant under the action of two groups:
  1. (1)
    The right action of the group $\mathrm{PSU}(2)$, which is the image of the unitary group $\mathrm{U}(2)$ under the projection $\mathrm{SL}_{2}(\mathbb{C})\to\mathrm{PSL}_{2}(\mathbb{C})$.
  2. (2)
    The left action of the group $G_{0}\leq\mathrm{PSL}_{2}(\mathbb{C})$ defined by:
    $G_{0}=\left\{\begin{bmatrix}x&iy\\ iz&w\end{bmatrix}\colon x,y,z,w\in\mathbb{R},\,xw+yz=1\right\}.$ (2)
In [GPPX22] the authors proved that, when restricted to the lattices $\mathrm{PSL}_{2}(\mathbb{Z})$ and $\mathrm{PSL}_{2}(\mathbb{Z}[i])$, this function defines an $L_{p}$-bounded Fourier multiplier for every $1<p<\infty$. They posed three related questions, namely:
  1. (1)
    Is this multiplier bounded in $L_{p}(\mathcal{L}\,\mathrm{PSL}_{2}(\mathbb{C}))$?
  2. (2)
    Is its restriction bounded in $L_{p}(\mathcal{L}\,\mathrm{PSL}_{2}(\mathbb{R}))$?
  3. (3)
    Are there more lattices $\Gamma\leq\mathrm{PSL}_{2}(\mathbb{C})$ for which the restriction of $m$ still defines a multiplier bounded in $L_{p}(\mathcal{L}\,\Gamma)$?
The two first questions are negatively answered by the work of Parcet, de la Salle and Tablate. Concretely, by [PdlST23, Corollary B2] and the fact that the Lie algebra of $\mathrm{PSL}_{2}(\mathbb{C})$ is simple (as a real Lie algebra) solves the problem.
In the present work we tackle the third question. Our main result concerns the family of groups $\Gamma_{n}=\mathrm{PSL}_{2}(\mathbb{Z}[\sqrt{-n}])$, and it can be stated follows:
Theorem A.
For any integer $n>0$, the symbol $m$ restricted to the group $\Gamma_{n}$ defines a bounded Fourier multiplier in $L_{p}(\mathcal{L}\,\Gamma_{n})$ for all $1<p<\infty$, whose norm satisfies:
$\|T_{m}\colon L_{p}(\mathcal{L}\,\Gamma_{n})\to L_{p}(\mathcal{L}\,\Gamma_{n})\|\lesssim\left(\frac{p^{2}}{p-1}\right)^{\beta},\quad\text{where }\beta=1+\log_{2}(1+\sqrt{2}).$
The proof consist in identifying a subgroup $K_{n}\leq\Gamma_{n}$ and a suitable character $\chi\colon K_{n}\to\mathbb{T}^{1}$ for which $m$ is left $(K_{n},\chi)$-invariant, and then proving by hand that (Cotlar) holds. This is a refinement of the argument in [GPPX22] for the case $n=1$, where the authors defined an auxiliary symbol $\widetilde{m}$ that is indeed $K_{1}$-invariant, and carried out the analogous computations that we present here in more generality.
Bianchi groups are another natural family of lattices in $\mathrm{PSL}_{2}(\mathbb{C})$ to consider. Indeed, for every square-free positive integer $n>0$, we define the $n$-th Bianchi group as $\Gamma_{n}^{\prime}=\mathrm{PSL}_{2}(\mathcal{O}_{-n})$, where $\mathcal{O}_{-n}$ denotes the ring of integers of the quadratic extension $\mathbb{Q}(\sqrt{-n})$. The explicit definition of $\Gamma_{n}^{\prime}$ depends on the class of $n$ modulus 4, since:
$\mathcal{O}_{-n}=\begin{cases}\mathbb{Z}[\sqrt{-n}]&\text{if }n\not\equiv-1\,(\operatorname{mod}4),\\ \mathbb{Z}\left[\frac{1+\sqrt{-n}}{2}\right]&\text{otherwise}.\end{cases}$
Therefore this family extends the one featuring in Theorem A. The problem is that the set where (Cotlar) fails is bigger in $\Gamma_{n}^{\prime}$ than in $\Gamma_{n}$. This set cannot be contained in a subgroup respect of which $m$ has some kind of invariance, which was a key step in our previous argument, and this is why the question of whether $m$ defines a bounded multiplier on $L_{p}(\mathcal{L}\,\Gamma_{n}^{\prime})$ is left open.
Acknowledgments
The author would like to thank his advisor, Adrián González-Pérez, for pointing him out the problem, as well as for his support and insightful conversations.
1.  Background
Group von Neumann algebras
Let $G$ be a discrete group and let $\lambda\colon G\to B(\ell_{2}(G))$ denote the left regular representation of $G$, that is, the unitary representation of $G$ asigning to every $g\in G$ the operator $\lambda_{g}\in B(\ell_{2}(G))$ given by $\lambda_{g}f(h)=f(g^{-1}h)$, for every $f\in\ell_{2}(G)$ and $h\in G$. The group von Neumann algebra of $G$, denoted here by $\mathcal{L}\,G$, is the operator algebra given by:
$\mathcal{L}\,G=\overline{\textrm{span}\{\lambda_{g}\colon g\in G\}}^{\textrm{WOT}},$
where closure is taken in the weak operator topology of $B(\ell_{2}(G))$. Notice that an arbitrary element $x\in\mathcal{L}\,G$ can be represented by a sum $x=\sum_{g\in G}x_{g}\lambda_{g}$, with $x_{g}\in\mathbb{C}$, converging in this topology.
The group von Neumann algebra $\mathcal{L}\,G$ comes equipped with a finite trace:
$\tau\colon\mathcal{L}\,G\to\mathbb{C},\quad x\mapsto\tau\left(\sum_{g\in G}x_{g}\lambda_{g}\right)=x_{e}.$
If $G$ is Abelian then $\mathcal{L}\,G$ is isomorphic (as von Neumann algebra) to $L_{\infty}(\widehat{G})$, where $\widehat{G}$ represents the dual group, and $\tau$ is the functional induced on $L_{\infty}(\widehat{G})$ by the Haar measure of $\widehat{G}$. In the non-commutative case, the trace $\tau$ above defined helps us to define $L_{p}$-spaces associated to $\mathcal{L}\,G$ without needing an underlying measure space. For a given $x\in\mathcal{L}\,G$ and $p\in[1,\infty]$ we define the norms:
$\|x\|_{p}=\tau(|x|^{p})^{1/p}\,\text{ if }1\leq p<\infty,\text{ and }\,\|x\|_{\infty}=\|x\|_{\mathcal{L}\,G}.$
The space $L_{p}(\mathcal{L}\,G)$ is defined as the closure of $B(\ell_{2}(G))$ respect to this norm. All of this can be done in more generality for non-discrete groups, using the Haar measure of $G$ and defining a weight $\tau$ instead of a trace, see [Ped79]. The $L_{p}$-spaces over von Neumann algebras can also be defined in more generality, see for example [PX03].
Non-commutative Fourier multipliers
A Fourier multiplier $T_{m}$ with symbol $m\colon G\to\mathbb{C}$ is an operator defined as:
$T_{m}\left(\sum_{g\in G}x_{g}\lambda_{g}\right)=\sum_{g\in G}m(g)x_{g}\lambda_{g},\quad\text{for }x=\sum_{g\in G}x_{g}\lambda_{g}\in\mathbb{C}G.$
Here $\mathbb{C}G$ denotes the space of elements with finite Fourier expansion. Notice that it is a dense subspace of every $L_{p}(\mathcal{L}\,G)$. If $T_{m}$ extends to a bounded operator $T_{m}\colon L_{p}(\mathcal{L}\,G)\to L_{p}(\mathcal{L}\,G)$, we say that $T_{m}$ is a bounded $L_{p}$-multiplier.
The study of general conditions for the symbol $m$ that ensure the $L_{p}$-boundedness of $T_{m}$ has been an active area of research both in the classical and the non-commutative case. As discussed in the Introduction, the key result we are going to use concerns the following version of Cotlar identity for non-commutative Fourier multipliers:
Theorem 1.1.
[GPPX22, Theorem A] Let $G$ be a locally compact unimodular group, $G_{0}\leq G$ an open subgroup and $\chi\colon G_{0}\to\mathbb{T}^{1}$ some character. Let $T_{m}$ be a Fourier multiplier whose symbol $m\colon G\to\mathbb{C}$ is bounded and left $(G_{0},\chi)$-invariant. If $m$ satisfies the identity:
$(m(g^{-1})-m(h))(m(gh)-m(g))=0,\text{ for all }g\in G\setminus G_{0}\text{ and}h\in G$
the $T_{m}$ is bounded in $L_{p}$ for all $1<p<\infty$. Moreover, its norm satisfies:
$\|T_{m}\colon L_{p}(\mathcal{L}\,G)\to L_{p}(\mathcal{L}\,G)\|\lesssim\left(\frac{p^{2}}{p-1}\right)^{\beta},\quad\text{with }\beta=\log_{2}(1+\sqrt{2}).$
Notice that if $G$ is discrete then any subgroup $G_{0}\leq G$ is open. The subgroup $G_{0}$ gives a range of flexibility to this result respect to the original one of Cotlar: taking a big subgroup $G_{0}$ increases the chances for the formula to hold, but makes it harder for $m$ to satisfy the invariance hypothesis. Also, the theorem holds even if we take $G_{0}$ an empty set, which allows us to recover the classical statement when $G=\mathbb{R}$.
2.  Description of the set where Cotlar identity fails
Let $m$ be the function defined in (1) and set $\Gamma_{n}=\mathrm{PSL}_{2}(\mathbb{Z}[\sqrt{-n}])$. As we shall prove later, our function ${\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{\Gamma_{n}}}$ is invariant (through a suitable character) respect to:
$K_{n}=\{g\in\Gamma_{n}\colon m(g)=0\},$ (2.1)
which turns out to be a subgroup of $\Gamma_{n}$. The goal of this section is to give an explicit description of this set. Along our proof, we will also give a description of the analogous set
$K_{n}^{\prime}=\{g\in\Gamma_{n}^{\prime}\colon m(g)=0\}$ (2.2)
[…]
The main theorem of this section (namely, Theorem B) allows us to decompose $K_{n}$ and $K_{n}^{\prime}$ as a combination of the four following disjoint sets:
$\begin{split}K_{n}^{+}&=\left\{\begin{bmatrix}x&y\sqrt{-n}\\ z\sqrt{-n}&w\end{bmatrix}\colon x,y,z,w\in\mathbb{Z},\;xw+nyz=1\right\},\\ K_{n}^{-}&=\left\{\begin{bmatrix}x\sqrt{-n}&y\\ z&w\sqrt{-n}\end{bmatrix}\colon x,y,z,w\in\mathbb{Z},\;nxw+yz=1\right\},\\ L_{n}^{+}&=\left\{\begin{bmatrix}a&-\overline{a}\\ c&\overline{c}\end{bmatrix}\colon a,c\in\mathcal{O}_{-n},\;\operatorname{Re}(a\overline{c})=\frac{1}{2}\right\},\quad\text{and}\\ L_{n}^{-}&=\left\{\begin{bmatrix}a&\overline{a}\\ c&-\overline{c}\end{bmatrix}\colon a,c\in\mathcal{O}_{-n},\;\operatorname{Re}(a\overline{c})=-\frac{1}{2}\right\}.\end{split}$ (2.3)
Theorem B.
[…]
  1. (1)
    $K_{n}=K_{n}^{+}\cup K_{n}^{-}$.
  2. (2)
[…]
We will divide the proof of the theorem in several lemmas. Concretely, Lemmas 2.4 and 2.6 combined prove the theorem for $K_{n}$, whereas Lemma 2.5 gives the extra information needed for $K_{n}^{\prime}$. Prior to that, we are going to make two remarks about the classification.
Remark 2.1.
The set $K_{n}$ is always a subgroup of $\Gamma_{n}$, but $K_{n}^{\prime}\subset\Gamma_{n}^{\prime}$ is not. This is because $K^{\prime}_{n}$ is not closed under products neither taking inverses. Let $l\in L_{n}^{+}$ and $g\in\mathrm{PSL}_{2}(\mathbb{C})$, then:
$m(lg)=\operatorname{sign}\left(|r_{1}(g)|-|r_{2}(g)|\right)$
where $r_{1}(g)$ and $r_{2}(g)$ denotes the first and second rows of $g$ as complex vectors in $\mathbb{C}^{2}$. Therefore $m(lg)\neq 0$ for many $g\in K_{n}^{\prime}$. On the other hand, notice that:
$(L_{n}^{+})^{-1}=\left\{\begin{bmatrix}c&a\\ -\overline{c}&\overline{a}\end{bmatrix}\colon\operatorname{Re}(a\overline{c})=1/2\right\},$
which is not contained in $K^{\prime}_{n}$ because of the classification we provided in Theorem B.
Remark 2.2.
The theorem does not apply for $K_{3}^{\prime}$. Notice that $\mathcal{O}_{-3}=\mathbb{Z}[\xi_{3}]$ where $\xi_{3}$ denotes a primitive $3$-root of the unit. A matrix as simple as:
$u=\begin{bmatrix}\xi_{3}&0\\ 0&\overline{\xi_{3}}\end{bmatrix}$
will be in $K_{3}^{\prime}$ but not in $K_{3}\cup L_{3}$. Also, since $m$ is right $\mathrm{PSU}(2)$-invariant, if we pick any $g\in K_{3}$ then $gu\in K_{3}^{\prime}$, but this product will not be in $K_{3}\cup L_{3}$ in general.
Lemma 2.3.
For any $g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in\mathrm{PSL}_{2}(\mathbb{C})$, it holds that $\operatorname{Re}(a\overline{c})\operatorname{Re}(b\overline{d})\geq-\frac{1}{4}$. Moreover, if $g\in\Gamma_{n}$, the right-hand side of the inequality can be improved to 00.
  • Proof.
    Suppose that $\operatorname{Re}(a\overline{c})\operatorname{Re}(b\overline{d})<0$. Then, by multiplying the equation $ad-bc=1$ by $\overline{cd}$ and taking real part we get that:
    $|\operatorname{Re}(b\overline{d})||c|^{2}+|\operatorname{Re}(a\overline{c})||d|^{2}=|\operatorname{Re}(c\overline{d})|\leq|c||d|.$
    Now we claim that any positive numbers $x,y,\alpha,\beta>0$ satisfying
    $\alpha x^{2}+\beta y^{2}\leq xy$ (2.4)
    must verify $\alpha\beta\leq\frac{1}{4}$. To prove the claim, just notice that (2.4) is equivalent to $\alpha u^{2}-u+\beta\leq 0$ with $u=x/y$, and this can only happen if the discriminant $1-4\alpha\beta$ is greater than or equal to 00.
    If $g\in\Gamma_{n}$, then both $\operatorname{Re}(a\overline{c})$ and $\operatorname{Re}(b\overline{d})$ are integers, so $\operatorname{Re}(a\overline{c})\operatorname{Re}(b\overline{d})$ must be indeed non-negative. ∎
Lemma 2.4.
For any $n\geq 0$, let $K_{n}$ and $K_{n}^{\prime}$ be the sets defined in (2.1) and (2.2) respectively. It holds that:
  1. (1)
    $K_{n}=\{g\in\Gamma_{n}\colon\operatorname{Re}(a\overline{c})=\operatorname{Re}(b\overline{d})=0\}$.
  2. (2)
[…]
  • Proof.
    For any $g\in\mathrm{PSL}_{2}(\mathbb{Z}[\sqrt{-n}])$, Lemma 2.3 says that $\operatorname{Re}(a\overline{c})\operatorname{Re}(b\overline{d})\geq 0$. Recall that $m(g)=\operatorname{Re}(a\overline{c})+\operatorname{Re}(b\overline{d})$, so $m(g)=0$ if and only if $\operatorname{Re}(a\overline{c})=\operatorname{Re}(b\overline{d})=0$. This proves the first point of the lemma. The second one follows from the same argument, but in this case $\operatorname{Re}(a\overline{c})$ and $\operatorname{Re}(b\overline{d})$ are demi-integers instead of integers since $\Gamma_{n}^{\prime}=\mathrm{PSL}_{2}(\mathbb{Z}[\frac{1+\sqrt{-n}}{2}])$. ∎
Lemma 2.5.
Let $g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}$ be an element of $\mathrm{PSL}_{2}(\mathbb{C})$.
  1. (1)
    $\operatorname{Re}(a\overline{c})=-\operatorname{Re}(b\overline{d})=\frac{1}{2}$ if and only if $c=\overline{d}$, $a=-\overline{b}$ and $\operatorname{Re}(a\overline{c})=\frac{1}{2}$.
  2. (2)
    $\operatorname{Re}(a\overline{c})=-\operatorname{Re}(b\overline{d})=-\frac{1}{2}$ if and only if $c=-\overline{d}$, $a=\overline{b}$ and $\operatorname{Re}(a\overline{c})=-\frac{1}{2}$.
  • Proof.
    In order to prove (1), first multiply the equation $ad-bc=1$ by $\overline{cd}$ and then take real part, resulting:
    $|d|^{2}+|c|^{2}=2\operatorname{Re}(\overline{cd}).$
    A basic algebraic computation shows that this equality is equivalent to:
    $|c-\overline{d}|^{2}=0,$
    This implies that $c=\overline{d}$. To obtain $a=-\overline{b}$, begin with the equation $\overline{ad}-\overline{bc}=0$, multiply it by $ab$ and take real part again, concluding the argument in the same fashion as before.
    Now, (2) follows from (1) just by considering $\tilde{g}=\begin{bmatrix}a&-b\\ -c&d\end{bmatrix}$ instead of $g$. ∎
Lemma 2.6.
Let $n\geq 0$ be an integer.
  1. (1)
    Let $g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in\Gamma_{n}$ verifying $\operatorname{Re}(a\overline{c})=\operatorname{Re}(b\overline{d})=0$. Then either $a,d\in i\mathbb{R}$ and $b,c\in\mathbb{R}$, or $a,d\in\mathbb{R}$ and $b,c\in i\mathbb{R}$.
  2. (2)
    Same holds if $g\in\Gamma_{n}^{\prime}$ for $n\equiv-1\,(\operatorname{mod}4)$, $n\neq 3$.
  • Proof.
    We’ll divide the argument in four cases, but first some notation: throughout this proof, for any given $x\in\mathbb{Z}[\sqrt{-n}]$ we will denote by $x_{1},x_{2}$ the integers satisfying $x=x_{1}+x_{2}\sqrt{-n}$. On the other hand, if $x\in\mathbb{Z}[\frac{1+\sqrt{-n}}{2}]$, then $x_{1},x_{2}$ will denote the integers satisfying $x=\frac{x_{1}+x_{2}\sqrt{-n}}{2}$.
    Case 1. Suppose that some entry is 00. We’ll make the proof for $a=0$ since all the cases are proven the same way. From $bc=-1$ it follows that $b=-c=\pm 1$, since these are the only units in the rings $\mathbb{Z}[\sqrt{-n}],\,\mathcal{O}_{-n}$, and therefore $d$ is purely imaginary.
    Case 2. Suppose that $g$ has non-zero entries and some entry is purely real or imaginary. From $\operatorname{Re}(a\overline{c})=\operatorname{Re}(b\overline{d})=0$ we know that there exist rational numbers $r,q\in\mathbb{Q}$ such that
    $c=r\sqrt{-n}a\quad\text{and}\quad d=q\sqrt{-n}b.$ (2.5)
    This also implies that:
    $ad-bc=(q-r)\sqrt{-n}ab=1,$ (2.6)
    and the statement follows from a simple checking.
    Case 3. Now we are proving by reductio ad absurdum that $g\in\Gamma_{n}$ have a purely real or purely imaginary entry. Suppose this is not the case, and also that $\operatorname{gcd}(a_{1},a_{2})=\operatorname{gcd}(b_{1},b_{2})=1$, since the general case follows in an analogous manner dividing first by the g.c.d. Let $r$ and $q$ be as in (2.5), and notice that (2.6) can be split into two real equations by taking real and imaginary part:
    $a_{1}b_{1}-na_{2}b_{2}=0,$ (2.7a)
    $(r-q)n(a_{1}b_{2}+b_{2}a_{1})=1.$ (2.7b)
    The definition of $r$ implies that its denominator, denoted by $\mathrm{den}(r)$, divides both $a_{1}$ and $na_{2}$. Therefore $\mathrm{den}(r)$ divides $n$. For the same reasons, $\mathrm{den}(q)$ also must divide $n$, so $(r-q)n$ is an integer.
    Because of (2.7b) we have that $a_{1}b_{2}+b_{2}a_{1}=\pm 1$. On the other hand, Equation (2.7a) implies that $a_{2}$ divides $b_{1}$ and $b_{2}$ divides $a_{1}$, so there exist integers $x,y$ such that $a_{1}=xb_{2}$, $b_{1}=ya_{2}$ and $xy=n$. All of this is summarized in the following two equations:
    $xy=n,$ (2.8a)
    $ya_{2}^{2}+xb_{2}^{2}=\pm 1.$ (2.8b)
    First equation says that $x$ and $y$ are non-zero and that they have the same sign, so $|ya_{2}^{2}+xb_{2}^{2}|\geq|a|^{2}+|b|^{2}$. However, both $|a|^{2}$ and $|b|^{2}$ are positive integers, so $|ya_{2}^{2}+xb_{2}^{2}|\geq 2$, which is a contradiction with the last equation.
    This proves (1). Now consider $g\in\Gamma_{n}^{\prime}$. The same argument as in Case 3 can be carried out, with the difference that now (2.8a) and (2.8b) have the form:
    $xy=n,$ (2.9)
    $|ya_{2}^{2}+xb_{2}^{2}|\leq 4.$
    Again the integers $x$ and $y$ have the same sign, so $4\geq|xb_{2}^{2}+ya_{2}^{2}|\geq|x|+|y|$. But any two integers $x,y$ satisfying this inequality verify $|xy|\leq 3$. Since $n>3$, we have a contradiction with the first equation, proving (2). ∎
3.  Proof of the Cotlar Identity
The sets $K_{n}^{+}$ and $K_{n}^{-}$ defined in (2.3) verify certain relations related to the invariance of $m$. Throughout this section, we set:
$\omega=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}\in\mathrm{PSL}_{2}(\mathbb{C}).$
It holds that $\omega K_{n}^{+}=K_{n}^{+}\omega=K_{n}^{-}.$ These identities, together with the fact that $K_{n}^{+}$ is a subgroup of $\Gamma_{n}$, implies easily that:
$K_{n}^{+}K_{n}^{-},\,K_{n}^{-}K_{n}^{+}\subset K_{n}^{-}\quad\text{and}\quad K_{n}^{-}K_{n}^{-}\subset K_{n}^{+}.$
We claim now that, because of these contentions, the function $\chi\colon K_{n}\to\mathbb{T}^{1}$ defined as:
$\chi(g)=\begin{cases}1&\text{if }g\in K_{n}^{+},\\ -1&\text{if }g\in K_{n}^{-},\end{cases}$
is a character. The following three lemmas prove that ${\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{\Gamma_{n}}}$ is left $(K_{n},\chi)$-invariant.
Lemma 3.1.
Let $g\in\mathrm{PSL}_{2}(\mathbb{C})$ and let $r_{1}(g)$ and $r_{2}(g)$ denote the first and second rows of $g$, respectively. There exist an unitary matrix $u\in\mathrm{PSU}(2)$ such that:
$g=\begin{bmatrix}s^{-1}&s^{-1}t\\ 0&s\end{bmatrix}u,$
with $s=|r_{2}(g)|$ and $t=\langle r_{1}(g),r_{2}(g)\rangle$, where the bracket represents the scalar product in $\mathbb{C}^{2}$.
  • Proof.
    This is just an explicit statement of the $\mathrm{ANK}$ decomposition for $\mathrm{PSL}_{2}(\mathbb{C})$. It can be proven directly as follows. Let $u$ be the (only) unitary matrix such that $r_{2}(g)u^{*}=(0,s)$ with $s>0$. Thus, $s=|r_{2}(g)|$. On the other hand, using that $\det(gu^{*})=1$, we get that $r_{1}(g)u^{*}=(s^{-1},\omega)$ for some $\omega\in\mathbb{C}$. This $\omega$ can be computed using that $\omega=s^{-1}\langle r_{1}(gu^{*}),r_{2}(gu^{*})\rangle=s^{-1}\langle r_{1}(g),r_{2}(g)\rangle$, which is the definition of $s^{-1}t$. ∎
Lemma 3.2.
For any $g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in\mathrm{PSL}_{2}(\mathbb{C})$, it holds that:
$\operatorname{Im}(b\overline{c}-a\overline{d})^{2}-4\operatorname{Re}(a\overline{c})\operatorname{Re}(b\overline{d})\leq 1.$
Moreover, if $g\in\Gamma_{n}$, then the right-hand side of the above inequality can be improved to 00.
  • Proof.
    Same computations as in the proof of [GPPX22, Lemma 5.3] shows that the left-hand side of the above expression can be written as $p(X)=-4X(1+X)$, where $X=na_{2}d_{2}+b_{1}c_{1}$. This proves the statement for $g\in\mathrm{PSL}_{2}(\mathbb{C})$. If $g\in\Gamma_{n}$, then $X$ is an integer and therefore $p(X)\in 4\mathbb{Z}$, which proves the second part of the statement. ∎
Lemma 3.3.
The symbol ${\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{\Gamma_{n}}}$ is right $K_{n}$-invariant and left $(K_{n},\chi)$-invariant.
  • Proof.
    It is immediate to check that $m(\omega g)=-m(g).$ On the other hand, $K_{n}^{+}$ and $K_{n}^{-}$ are contained respectively in $G_{0}$ and $\omega G_{0}$, where $G_{0}$ is the group defined in the Introduction by (2). Since $m$ is invariant by the left action of $G_{0}$, it follows that ${\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{\Gamma_{n}}}$ is left $(K_{n},\chi)$-invariant.
    For the right invariance, let’s take $g\in\Gamma_{n}$ and $h\in K_{n}$. If $g\in K_{n}$, it is immediate that $m(gh)=m(g)=0$, so we rule out this case. Let’s write $g$ and $h$ as
    $g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\quad\text{and}\quad h=\begin{bmatrix}s^{-1}&s^{-1}t\\ 0&s\end{bmatrix}u,$
[…]
    $\begin{split}\operatorname{Re}\langle r_{1}(gh),r_{2}(gh)\rangle&=\operatorname{Re}(a\overline{c})(1+(\operatorname{Im}t)^{2})s^{-2}+\operatorname{Re}(b\overline{d})s^{2}+\operatorname{Im}(b\overline{c}-a\overline{d})\operatorname{Im}t\\ &=\begin{bmatrix}s^{-1}\operatorname{Im}t&s\end{bmatrix}\begin{bmatrix}2\operatorname{Re}(a\overline{c})&\operatorname{Im}(b\overline{c}-a\overline{d})\\ \operatorname{Im}(b\overline{c}-a\overline{d})&2\operatorname{Re}(b\overline{d})\end{bmatrix}\begin{bmatrix}s^{-1}\operatorname{Im}t\\ s\end{bmatrix}+s^{-2}\operatorname{Re}(a\overline{c})\end{split}$ (3.1)
    The Lemma 3.2 says that the determinant of the matrix in (3.1) is always non-negative. Therefore, this matrix will be semidefinite positive if $\operatorname{Re}(a\overline{c})\geq 0$ and $\operatorname{Re}(b\overline{d})\geq 0$ and semidefinte negative otherwise. In both cases, it implies that $m(gh)m(g)\geq 0$. Since $gh\not\in K_{n}$, it follows that $m(gh)=m(g)$, proving the statement. ∎
Lemma 3.4.
Let $g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in\Gamma_{n}$. Then
$m(g)m(g^{t})\operatorname{Re}(a\overline{d}+b\overline{c})\geq 0,$
where $g^{t}$ denotes the transpose of $g$.
  • Proof.
    If $g$ or $g^{t}$ are in $K_{n}$, the result is immediate. If they are not, we know $m(g)$ has the same sign as $\operatorname{Re}(a\overline{c})$ or $\operatorname{Re}(b\overline{d})$, depending on which one is non-zero. We’ll suppose that both $\operatorname{Re}(a\overline{c})$ and $\operatorname{Re}(a\overline{b})$ are non-zero, since the rest of the cases comes from applying this one to $\omega g$, $g\omega$ or $\omega g\omega$.
    Under this hypothesis, the statement is equivalent to
    $\operatorname{Re}(a\overline{c})\operatorname{Re}(a\overline{d})\operatorname{Re}(a\overline{d}+b\overline{c})\geq 0.$
    From the proof of [GPPX22, Proposition 5.8] we know that the left-hand side of the inequality equals $p(X)=(AX+B)(2X+1)$ with $A=n(a_{1}^{2}+a_{2}^{2})$, $B=na_{2}^{2}$ and $X=b_{1}c_{1}+na_{2}d_{2}$. Since $X$ is an integer and the roots of the polynomial $p$ have modulus lesser or equal than $1$, we conclude the statement. ∎
  • Proof of Theorem A.
    We are going to prove that the symbol ${\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{\Gamma_{n}}}$ satisfies (Cotlar) relative to $K_{n}$, that is:
    $(m(g^{-1})-m(h))(m(gh)-m(g))=0,\text{ for all }g\in\Gamma_{n}\setminus K_{n}\text{ and }h\in\Gamma_{n}.$
    If $h\in K_{n}$, the equality follows from the right $K_{n}$-invariance of $m$ proven in Lemma 3.3. Now, suppose that $h\not\in K_{n}$ and $m(g^{-1})\neq m(h)$. We have to prove that $m(gh)=m(g)$. Since the hypothesis $m(g^{-1})\neq m(h)$ implies that $gh\not\in K_{n}$, it suffies to prove that $m(gh)m(g)\geq 0$. We write:
    $g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\quad\text{and}\quad h=\begin{bmatrix}s^{-1}&s^{-1}t\\ 0&s\end{bmatrix}u,$
    using Lemma 3.1 to decompose $h$ into an upper-triangular matrix and an unitary one. Now, a computation shows that:
    $\begin{split}m(gh)m(g)&=\operatorname{sign}\Big{(}\operatorname{Re}(a\overline{c}+b\overline{d})\operatorname{Re}(a\overline{c})s^{-2}(1+(\operatorname{Re}t)^{2})\\ &\qquad+\operatorname{Re}(a\overline{c}+b\overline{d})\operatorname{Re}(a\overline{d}+b\overline{c})\operatorname{Re}t\\ &\qquad+\operatorname{Re}(a\overline{c}+b\overline{d})\big{[}\operatorname{Re}(a\overline{c})s^{-2}(\operatorname{Im}t)^{2}+\operatorname{Re}(b\overline{d})s^{2}+\operatorname{Im}(b\overline{c}-a\overline{d})\operatorname{Im}t\big{]}\Big{)}\\ &=\operatorname{sign}\Big{(}(\mathrm{I})+(\mathrm{II})+(\mathrm{III})\Big{)}\end{split}$
    Now notice that $(\mathrm{I})$ is non-negative because of Lemma 2.3 and the fact that $s>0$. Also, $(\mathrm{II})$ is non-negative because $\operatorname{Re}t$ has the same sign as $m(h)=-m(g^{-1})=m(g^{t})$, so we can apply Lemma 3.4. Finally, $(\mathrm{III})$ is non-negative because of Lemma 3.2, which implies that each factor of the product has the same sign as $m(g)$ or is zero. ∎
Remark 3.5.
We still don’t know if the Fourier multiplier given by ${\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{\Gamma_{n}^{\prime}}}$ is bounded or not in $L_{p}(\mathcal{L}\,\Gamma_{n}^{\prime})$, but what can be proven is that this symbol do not verify a Cotlar identity as in Theorem 1.1 respect to any possible subgroup of $\Gamma_{n}^{\prime}$. To see this, suppose that ${\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{\Gamma_{n}^{\prime}}}$ is $(G_{0},\chi)$-invariant for some subgroup $G_{0}\leq\Gamma_{n}^{\prime}$ and some character $\chi$ on $G_{0}$. We claim that $G_{0}\cap L_{n}^{+}=\varnothing$. Firstly, let $a\colon\Gamma_{n}^{\prime}\to\Gamma_{n}^{\prime}$ the map that permutes the two rows of a matrix and multiplies the first column by $-1$. If $G_{0}\cap L_{n}^{+}\neq\varnothing$, by the formula featuring in Remark 2.1 it would hold that for any $h\in\Gamma_{n}^{\prime}$ and any $l\in G_{0}\cap L_{n}^{+}$:
$m(lh)=\chi(l)m(h)=\chi(l)m(a(h))=m(la(h))=-m(lh),$
which is of course impossible. On the other hand, fix an $l\in L^{+}_{n}$ whose inverse is not in $K_{n}^{\prime}$. Then in order to Cotlar identity to hold, one needs that:
$m(l^{-1})=m(h),\quad\text{for any }h\in\Gamma_{n}^{\prime}\text{ such that }m(lh)\neq 0.$
Pick $h\in\Gamma_{n}^{\prime}$ any element which verifies this equation. Let $h^{\prime}$ be given by $h^{\prime}=\omega h\omega$, where $\omega$ is the matrix defined at the beginning of this section. Notice that $m(lh^{\prime})=-m(lh)\neq 0$, but $m(h^{\prime})=-m(h)$. Therefore Cotlar identity must fail when applied to $l$ and $h^{\prime}$.