### v2

# A note on Hilbert transform over lattices of $\mathrm{PSL}_{2}(\mathbb{C})$

Abstract.
González-Pérez, Parcet and Xia introduced recently a framework to study $L_{p}$-boundedness of certain families of idempotent multipliers on von Neumann algebras. It includes symbols $m\colon\mathrm{PSL}_{2}(\mathbb{C})\to\mathbb{R}$ arising from lifting the indicator function of a partition $\{\Sigma^{+},\Sigma^{+},\Sigma^{-}\}$ of the hyperbolic space $\mathbb{H}^{3}$ to its isometry group $\mathrm{PSL}_{2}(\mathbb{C})$. The boundedness of $T_{m}$ on $L_{p}(\mathcal{L}\mathrm{PSL}_{2}(\mathbb{C}))$ was disproved by Parcet, de la Salle and Tablate. Nevertheless, we will show that this Fourier multiplier is bounded when restricted to the arithmetic lattices $\mathrm{PSL}_{2}(\mathbb{Z}[\sqrt{-n}])$, solving a question left open by the first named authors.
Key words and phrases:
Group von Neumann algebras, Hilbert transform, Non-commutative Lp spaces, Non-commutative harmonic analysis
1991 Mathematics Subject Classification:
43A22, 46L52
Introduction
The boundedness problem for Fourier multipliers on $L_{p}$-spaces has always played a central role in harmonic analysis. One of the most studied examples is the Hilbert transform, defined as $\widehat{Hf}(\xi)=i\,\mathrm{sign}(\xi)\widehat{f}(\xi)$ for $f\in L_{2}(\mathbb{R})$. Although $H$ was already known to be bounded in $L_{p}(\mathbb{R})$ for $1<p<\infty$, in 1955 Cotlar [3] gave a very simple proof of this fact using the following identity:
$|Hf|^{2}=2H(f\cdot Hf)-H(H(|f|^{2})).$ (Classical Cotlar)
This is known nowadays as the Cotlar identity. His proof uses that $H$ is bounded in $L_{2}(\mathbb{R})$ and that, by a recursive use of (Classical Cotlar), it also must be bounded in every $p=2^{k}$ for $k\geq 1$. Interpolation and the fact that $H$ is self-adjoint complete the proof.
Mei and Ricard [5] introduced the Cotlar identity in the non-commutative setting in order to study Hilbert transforms over free groups and amalgamated free products of von Neumann algebras. In the recent work of González-Pérez, Parcet and Xia [4] the authors developed a systematic approach to study Cotlar identities for Fourier multipliers in non-Abelian groups. Let $G$ be an unimodular group, $\mathcal{L}G$ the von Neumann algebra of $G$ and $G_{0}\subset G$ an open subgroup. Consider $m\colon G\to\mathbb{C}$ a symbol on $G$ and $T_{m}$ the corresponding Fourier multiplier on $\mathcal{L}G$. Then the formula:
$(m(g^{-1})-m(h))(m(gh)-m(g))=0,\quad\text{for all }g\in G\setminus G_{0},\,h\in G,$ (Cotlar)
is a translation of (Classical Cotlar) for $T_{m}$ in terms of its symbol. The main result in [4] states that any $m$ which is bounded, left $G_{0}$-invariant and verifies (Cotlar) defines a bounded multiplier in $L_{p}(\mathcal{L}\,G)$ for all $1<p<\infty$.
The subgroup $G_{0}$ represents a set in which the Cotlar identity may fail. In the argument, this failure is balanced by the invariance of $m$ with respect to $G_{0}$. Therefore this formulation of the theorem allows more flexibility in terms of the multiplier than the original one. However, the hypothesis of invariance can be relaxed even further. If $\chi\colon G_{0}\to\mathbb{T}^{1}$ is a character, it is enough for the result to hold that $m$ verifies:
$m(gh)=\chi(g)m(h)\quad\text{for all }g\in G_{0},\,h\in G.$
We say in this case that $m$ is left $(G_{0},\chi)$-equivariant, and of course the $G_{0}$-invariance is recovered when $\chi$ is the trivial character.
Hilbert transform in $\mathrm{PSL}_{2}(\mathbb{C})$
Recall that $\mathrm{PSL}_{2}(\mathbb{C})$, which is the quotient of the $2\times 2$ complex matrices with determinant $1$ by its center, can be identified with the group of orientation-preserving isometries of the three dimensional hyperbolic space $\mathbb{H}^{3}$. This identification can be made explicit in various ways. Here we give one using the upper-space model of $\mathbb{H}^{3}$ and quaternions. Let $i,j,k$ denote the usual three quaternionic units, and let’s define:
$\mathbb{H}^{3}=\{x+yi+rj\colon x,y,r\in\mathbb{R},\,r>0\}.$
Doing so, $\mathbb{H}^{3}$ is exactly the subspace $\mathbb{C}+\mathbb{R}_{>0}j$ of the quaternions. Now, for a given $\omega\in\mathbb{H}^{3}$ we set:
$g\cdot\omega=(a\omega+b)(c\omega+d)^{-1},\quad\text{for }g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in\mathrm{PSL}_{2}(\mathbb{C}).$
It is possible to compute the inverse of a quaternion using its conjugate and modulus. This leads to a more explicit formula for the action of $g\in\mathrm{PSL}_{2}(\mathbb{C})$ on the element $\omega=z+rj\in\mathbb{C}+\mathbb{R}_{>0}j$, namely:
$g\cdot\omega=\frac{a\overline{c}|z+rj|^{2}+b\overline{d}+a\overline{d}z+b\overline{c}\overline{z}+rj}{|c(z+rj)+d|^{2}}$
This is a well-defined action of $\mathrm{PSL}_{2}(\mathbb{C})$ on $\mathbb{H}^{3}$. Indeed, $\mathrm{PSL}_{2}(\mathbb{C})$ acts by orientation-preserving isometries on $\mathbb{H}^{3}$ when equipped with the usual Riemannian metric:
$ds^{2}=\frac{dx^{2}+dy^{2}+dr^{2}}{r^{2}},$
and it is the full group of such isometries (see [1] for more details).
On the other hand, a group $G$ acting on a set $X$ induces a multiplier on $G$ as follows: first choose a point $x_{0}\in X$ and two disjoint subsets $X^{+},X^{-}\subset X$. Let $m$ be the map $m\colon G\to\mathbb{C}$ defined for each $g\in G$ as:
$m(g)=\begin{cases}1&\text{if }g\cdot x_{0}\in X^{+},\\ -1&\text{if }g\cdot x_{0}\in X^{-},\\ 0&\text{otherwise}.\end{cases}$
Even if the final multiplier depends on $x_{0}$ and also on the partition given by $X^{+}$ and $X^{-}$, the boundedness of the multiplier is preserved by changing $x_{0}$ for any other point in the same $G$-orbit or using the sets $\{g\cdot X^{+},g\cdot X^{-}\}$, with $g\in G$, instead of $\{X^{+},X^{-}\}$. Back to the action of $\mathrm{PSL}_{2}(\mathbb{C})$ on the hyperbolic space, we are choosing the base point in $\mathbb{H}^{3}$ given by $j$ in our quaternionic parametrization, and the following partition:
$\Sigma^{+}=\{\omega\in\mathbb{H}^{3}\colon\mathrm{Re}(\omega)>0\},\quad\text{and}\quad\Sigma^{-}=\{\omega\in\mathbb{H}^{3}\colon\mathrm{Re}(\omega)<0\}.$
This procedure induces a symbol $m$ in $\mathrm{PSL}_{2}(\mathbb{C})$ that is explicitly given by:
$m(g)=\mathrm{sign}(\mathrm{Re}(a\overline{c}+b\overline{d})),\quad\text{with }g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in\mathrm{PSL}_{2}(\mathbb{C}).$ (1)
The dividing frontier $\Sigma=\mathbb{H}^{3}\smallsetminus(\Sigma^{+}\cup\Sigma^{-})=\{\omega\in\mathbb{H}^{3}\colon\mathrm{Re}(\omega)=0\}$ is a hyperbolic plane, which determines the symbol $m$ up to a sign. Since the action of $\mathrm{PSL}_{2}(\mathbb{C})$ is transitive both on points and hyperbolic planes in $\mathbb{H}^{3}$, the boundedness of the multiplier defined by $m$ on $L_{p}(\mathcal{L}\,\mathrm{PSL}_{2}(\mathbb{C}))$ will remain the same under any other choice of the kind. Also, it is worth noticing that $m$ is easily shown invariant under the action of two groups:
  1. i.
    The right action of the group $\mathrm{PSU}(2)$, which is the image of the unitary group $\mathrm{U}(2)$ under the projection $\mathrm{SL}_{2}(\mathbb{C})\to\mathrm{PSL}_{2}(\mathbb{C})$.
  2. ii.
    The left action of the group $G_{0}\leq\mathrm{PSL}_{2}(\mathbb{C})$ defined by:
    $G_{0}=\left\{\begin{bmatrix}x&iy\\ iz&w\end{bmatrix}\colon x,y,z,w\in\mathbb{R},\,xw+yz=1\right\}.$ (2)
In [4] the authors proved that, when restricted to the lattices $\mathrm{PSL}_{2}(\mathbb{Z})$ and $\mathrm{PSL}_{2}(\mathbb{Z}[i])$, this function defines an $L_{p}$-bounded Fourier multiplier for every $1<p<\infty$. They posed three related questions, namely:
  1. i.
    Is this multiplier bounded in $L_{p}(\mathcal{L}\,\mathrm{PSL}_{2}(\mathbb{C}))$?
  2. ii.
    Is its restriction bounded in $L_{p}(\mathcal{L}\,\mathrm{PSL}_{2}(\mathbb{R}))$?
  3. iii.
    Are there more lattices $\Gamma\leq\mathrm{PSL}_{2}(\mathbb{C})$ for which the restriction of $m$ still defines a multiplier bounded in $L_{p}(\mathcal{L}\,\Gamma)$?
The two first questions are negatively answered by the work of Parcet, de la Salle and Tablate. Concretely, by [6, Corollary B2] and the fact that the Lie algebra of $\mathrm{PSL}_{2}(\mathbb{C})$ is simple (as a real Lie algebra) solves the problem.
In the present work we tackle the third question. Our main result concerns the family of groups $\Gamma_{n}=\mathrm{PSL}_{2}(\mathbb{Z}[\sqrt{-n}])$, and it can be stated follows:
Theorem A.
For any integer $n>0$, the symbol $m$ restricted to the group $\Gamma_{n}$ defines a bounded Fourier multiplier in $L_{p}(\mathcal{L}\,\Gamma_{n})$ for all $1<p<\infty$, whose norm satisfies:
$\|T_{m}\colon L_{p}(\mathcal{L}\,\Gamma_{n})\to L_{p}(\mathcal{L}\,\Gamma_{n})\|\lesssim\left(\frac{p^{2}}{p-1}\right)^{\beta},\quad\text{where }\beta=1+\log_{2}(1+\sqrt{2}).$
The proof consist in identifying a subgroup $K_{n}\leq\Gamma_{n}$ and a suitable character $\chi\colon K_{n}\to\mathbb{T}^{1}$ for which $m$ is left $(K_{n},\chi)$-equivariant, and then proving by hand that (Cotlar) holds. Using results and ideas from [9], we refined the argument in [4] for the case $n=1$, where the authors defined an auxiliary symbol $\widetilde{m}$ that is indeed $K_{1}$-invariant, and carried out the analogous computations that we present here in more generality.
Bianchi groups are another natural family of lattices in $\mathrm{PSL}_{2}(\mathbb{C})$ to consider, which were introduced by Bianchi in [2] as a generalization of the group $\mathrm{PSL}_{2}(\mathbb{Z})$. For every square-free positive integer $n>0$, we define the $n$-th Bianchi group as $\Gamma_{n}^{\prime}=\mathrm{PSL}_{2}(\mathcal{O}_{-n})$, where $\mathcal{O}_{-n}$ denotes the ring of integers of the quadratic extension $\mathbb{Q}(\sqrt{-n})$. The explicit definition of $\Gamma_{n}^{\prime}$ depends on the class of $n$ modulus 4, since:
$\mathcal{O}_{-n}=\begin{cases}\mathbb{Z}[\sqrt{-n}]&\text{if }n\not\equiv-1\,(\operatorname{mod}4),\\ \mathbb{Z}\left[\frac{1+\sqrt{-n}}{2}\right]&\text{otherwise}.\end{cases}$
Therefore this family extends the one featuring in Theorem A when $n\equiv-1\,(\operatorname{mod}4)$. In this case, the problem is that the set where (Cotlar) fails is bigger in $\Gamma_{n}^{\prime}$ than in $\Gamma_{n}$. This set cannot be contained in a subgroup with respect to which $m$ has some kind of invariance, and this is why the Cotlar identity cannot hold in every Bianchi group $\Gamma_{n}^{\prime}$ with $n\equiv-1\,(\operatorname{mod}4)$.
The question of whether $m$ defines a bounded multiplier on $L_{p}(\mathcal{L}\,\Gamma_{n}^{\prime})$ is left open in this case, but we are still able to prove that, choosing a different hyperbolic plane $\Sigma$ to induce our multiplier, one can still get symbols that satisfy Cotlar identity in most of Bianchi groups (all but $n=3$). Moreover, this approach also allows us to prove the Cotlar identity for the original $m$ in $\mathrm{PSL}_{2}(\mathbb{Z}[i])$ in a much simpler way than any of the previous proofs.
1. Background
Group von Neumann algebras
Let $G$ be a discrete group and let $\lambda\colon G\to B(\ell_{2}(G))$ denote the left regular representation of $G$, that is, the unitary representation of $G$ assigning to every $g\in G$ the operator $\lambda_{g}\in B(\ell_{2}(G))$ given by $\lambda_{g}f(h)=f(g^{-1}h)$, for every $f\in\ell_{2}(G)$ and $h\in G$. The group von Neumann algebra of $G$, denoted here by $\mathcal{L}\,G$, is the operator algebra given by:
$\mathcal{L}\,G=\overline{\textrm{span}\{\lambda_{g}\colon g\in G\}}^{\textrm{WOT}},$
where closure is taken in the weak operator topology of $B(\ell_{2}(G))$. Notice that an arbitrary element $x\in\mathcal{L}\,G$ can be represented by a sum $x=\sum_{g\in G}x_{g}\lambda_{g}$, with $x_{g}\in\mathbb{C}$.
The group von Neumann algebra $\mathcal{L}\,G$ comes equipped with a finite trace:
$\tau\colon\mathcal{L}\,G\to\mathbb{C},\quad x\mapsto\tau\left(\sum_{g\in G}x_{g}\lambda_{g}\right)=x_{e}.$
If $G$ is Abelian then $\mathcal{L}\,G$ is isomorphic (as von Neumann algebra) to $L_{\infty}(\widehat{G})$, where $\widehat{G}$ represents the dual group, and $\tau$ is the functional induced on $L_{\infty}(\widehat{G})$ by the Haar measure of $\widehat{G}$. In the non-commutative case, the trace $\tau$ above defined helps us to define $L_{p}$-spaces associated to $\mathcal{L}\,G$ without needing an underlying measure space. For a given $x\in\mathcal{L}\,G$ and $p\in[1,\infty]$ we define the norms:
$\|x\|_{p}=\tau(|x|^{p})^{1/p}\,\text{ if }1\leq p<\infty,\text{ and }\,\|x\|_{\infty}=\|x\|_{\mathcal{L}\,G}.$
The space $L_{p}(\mathcal{L}\,G)$ is defined as the completion of $B(\ell_{2}(G))$ with respect to this norm. All of this can be done in more generality for non-discrete groups, using the Haar measure of $G$ and defining a weight $\tau$ instead of a trace, see [7]. The $L_{p}$-spaces over von Neumann algebras can also be defined in more generality, see for example [8].
Non-commutative Fourier multipliers
A Fourier multiplier $T_{m}$ with symbol $m\colon G\to\mathbb{C}$ is an operator defined as:
$T_{m}\left(\sum_{g\in G}x_{g}\lambda_{g}\right)=\sum_{g\in G}m(g)x_{g}\lambda_{g},\quad\text{for }x=\sum_{g\in G}x_{g}\lambda_{g}\in\mathbb{C}G.$
Here $\mathbb{C}G$ denotes the space of elements with finite Fourier expansion. Notice that it is a dense subspace of every $L_{p}(\mathcal{L}\,G)$ for $1\leq p<\infty$. If $T_{m}$ extends to a bounded operator $T_{m}\colon L_{p}(\mathcal{L}\,G)\to L_{p}(\mathcal{L}\,G)$, we say that $T_{m}$ is a bounded $L_{p}$-multiplier.
The study of general conditions for the symbol $m$ that ensure the $L_{p}$-boundedness of $T_{m}$ has been an active area of research both in the classical and the non-commutative case. As discussed in the Introduction, the key result we are going to use concerns the following version of Cotlar identity for non-commutative Fourier multipliers:
Theorem 1.1.
[4, Theorem A] Let $G$ be a discrete group, $G_{0}\leq G$ a subgroup and $\chi\colon G_{0}\to\mathbb{T}^{1}$ some character. Let $T_{m}$ be a Fourier multiplier whose symbol $m\colon G\to\mathbb{C}$ is bounded and left $(G_{0},\chi)$-equivariant. If $m$ satisfies the identity:
$(m(g^{-1})-m(h))(m(gh)-m(g))=0,\text{ for all }g\in G\setminus G_{0}\text{ and}h\in G$
the $T_{m}$ is bounded in $L_{p}$ for all $1<p<\infty$. Moreover, its norm satisfies:
$\|T_{m}\colon L_{p}(\mathcal{L}\,G)\to L_{p}(\mathcal{L}\,G)\|\lesssim\left(\frac{p^{2}}{p-1}\right)^{\beta},\quad\text{with }\beta=\log_{2}(1+\sqrt{2}).$
The subgroup $G_{0}$ gives a range of flexibility to this result with respect to the original one of Cotlar: taking a big subgroup $G_{0}$ increases the chances for the formula to hold, but makes it harder for $m$ to satisfy the invariance hypothesis.
Hyperbolic planes, their boundaries and Möebius transformations
As we said, the group $\mathrm{PSL}_{2}(\mathbb{C})$ acts transitively on the set of pairs $(p,\Sigma)$ where $\Sigma$ is an hyperbolic plane embedded in $\mathbb{H}^{3}$ and $p$ is a point in $\Sigma$. When working with the upper-half space model $\mathbb{H}^{3}=\mathbb{C}+\mathbb{R}_{>0}j$, hyperbolic planes can be identified with half-planes and semispheres perpendicular (in the Euclidean sense) to $\mathbb{C}$. This induces a bijection between the set of hyperbolic planes in $\mathbb{H}^{3}$ and the set of generalized circles in ${}^\widehat{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$, that is, the set of lines and circles in $\mathbb{C}$. We will denote by $\partial\Sigma$ the generalized circle associated to $\Sigma$ by this correspondence, and we will call it the boundary of $\Sigma$.
Notice also that, for any given hyperplane $\Sigma$ and any $g\in\mathrm{PSL}_{2}(\mathbb{C})$, it holds that $\partial(g\cdot\Sigma)=g\cdot\partial\Sigma$, where $g$ is acting by Möbius transformation on the right-hand side. The action of Bianchi groups by Möbius transformations have been extensively studied in [9]. We will introduce now several results and concepts in that article, that we will make use of.
Let ${}^\widehat{\mathbb{R}}=\mathbb{R}\cup\{\infty\}$ be the extended real line. The Bianchi group $\Gamma_{n}^{\prime}$ acts on $\widehat{\mathbb{R}}$ in a controlled way:
Proposition 1.2.
[9, Proposition 4.4] If $n\neq 3$ and $g\in\Gamma_{n}^{\prime}$, then $\widehat{\mathbb{R}}$ and $g\cdot\widehat{\mathbb{R}}$ may only intersect tangentially.
On the other hand, for a given $g\in\mathrm{PGL}_{2}(\mathbb{C})$ with $|\mathrm{det}(g)|=1$, the quantities:
$\alpha=i(a\overline{d}-b\overline{c}),\quad\beta=-2\mathrm{Im}(c\overline{d}),\quad\beta^{\prime}=-2\mathrm{Im}(a\overline{b}),\quad\text{where }g=\begin{bmatrix}a&b\\ c&d\end{bmatrix},$ (1)
describe the image of $\widehat{\mathbb{R}}$ by $g$ in the following way:
$g\cdot\widehat{\mathbb{R}}=\left\{X/Y\in\widehat{\mathbb{C}}\colon\beta X\overline{X}-\alpha Y\overline{X}-\overline{\alpha}X\overline{Y}+\beta^{\prime}Y\overline{Y}=0\right\}.$
Proposition 1.3.
[9, Propositions 3.5 and 3.7] The coefficients $\alpha,$ $\beta$ and $\beta^{\prime}$ defined as above verify that:
  1. i.
    $\beta\beta^{\prime}=|\alpha|^{2}-1$,
  2. ii.
    the generalized circle $g\cdot\widehat{\mathbb{R}}$ goes through 00 if and only if $b^{\prime}=0$,
  3. iii.
    the generalized circle $g\cdot\widehat{\mathbb{R}}$ is indeed a line if and only if $b=0$,
  4. iv.
    if $b=0$, then $\alpha$ is a unit vector perpendicular to $g\cdot\widehat{\mathbb{R}}$,
  5. v.
    if $b\neq 0$, then $g\cdot\widehat{\mathbb{R}}$ is a circle of center $\alpha/\beta$ and radius $1/|\beta|$.
2. Description of the set where Cotlar identity fails
Throughout the rest of the paper, we will denote by $\tau$ and $\tau^{\prime}$ the following matrices:
$\tau=\begin{bmatrix}i&0\\ 0&1\end{bmatrix},\quad\text{and}\quad\tau^{\prime}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}.$
Let $m$ be the function defined in (1) and set $\Gamma_{n}=\mathrm{PSL}_{2}(\mathbb{Z}[\sqrt{-n}])$. As we shall prove later, our function ${\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{\Gamma_{n}}}$ is invariant (through a suitable character) with respect to:
$K_{n}=\{g\in\Gamma_{n}\colon m(g)=0\},$ (2)
which turns out to be a subgroup of $\Gamma_{n}$. The goal of this section is to give an explicit description of this set. Along our proof, we will also give a description of the analogous set
$K_{n}^{\prime}=\{g\in\Gamma_{n}^{\prime}\colon m(g)=0\}$ (3)
[…]
The main theorem of this section (namely, Theorem B) allows us to decompose $K_{n}$ and $K_{n}^{\prime}$ as a combination of the four following disjoint sets:
$\begin{split}K_{n}^{+}&=\left\{\begin{bmatrix}x&y\sqrt{-n}\\ z\sqrt{-n}&w\end{bmatrix}\colon x,y,z,w\in\mathbb{Z},\;xw+nyz=1\right\},\\ K_{n}^{-}&=\left\{\begin{bmatrix}x\sqrt{-n}&y\\ z&w\sqrt{-n}\end{bmatrix}\colon x,y,z,w\in\mathbb{Z},\;nxw+yz=-1\right\},\\ L_{n}^{+}&=\left\{\begin{bmatrix}a&-\overline{a}\\ c&\overline{c}\end{bmatrix}\colon a,c\in\mathcal{O}_{-n},\;\mathrm{Re}(a\overline{c})=\frac{1}{2}\right\},\quad\text{and}\\ L_{n}^{-}&=\left\{\begin{bmatrix}a&\overline{a}\\ c&-\overline{c}\end{bmatrix}\colon a,c\in\mathcal{O}_{-n},\;\mathrm{Re}(a\overline{c})=-\frac{1}{2}\right\}.\end{split}$ (4)
Lemma 2.1.
Let $g\in\mathrm{PSL}_{2}(\mathbb{C})$, $G_{0}\leq\mathrm{PSL}_{2}(\mathbb{C})$ be the group defined in (2) and:
[…]
Then it holds that:
  1. i.
    $g\cdot i\widehat{\mathbb{R}}=i\widehat{\mathbb{R}}$ if and only if $g\in G_{0}\cup\tau^{\prime}G_{0}$,
  2. ii.
    $g\cdot\mathbb{S}^{1}=i\widehat{\mathbb{R}}$ if and only if $g\in L\cup\tau^{\prime}L$.
Proof.
Notice that $g\cdot i\widehat{\mathbb{R}}=i\widehat{\mathbb{R}}$ if and only if $\sigma(g)\cdot\widehat{\mathbb{R}}=\widehat{\mathbb{R}}$. It is well-known that:
$\{g\in\mathrm{PSL}_{2}(\mathbb{C})\colon g\cdot\widehat{\mathbb{R}}=\widehat{\mathbb{R}}\}=\mathrm{PSL}_{2}(\mathbb{R})\cup\begin{bmatrix}i&0\\ 0&-i\end{bmatrix}\mathrm{PSL}_{2}(\mathbb{R}).$
The first point of the statement follows immediately.
We claim now that any $g\in L$ verifies $g\cdot\mathbb{S}^{1}=i\widehat{\mathbb{R}}$. This is because $g^{-1}\cdot 0,\,g^{-1}\cdot\infty$ and $g^{-1}\cdot i$ can be very easily checked to be all complex numbers in $\mathbb{S}^{1}$. On the other hand, given any $g_{0}\in\mathrm{PSL}_{2}(\mathbb{C})$ such that $g_{0}\cdot\mathbb{S}^{1}=i\widehat{\mathbb{R}}$, the set of $g\in\mathrm{PSL}_{2}(\mathbb{C})$ satisfying $g\cdot\mathbb{S}^{1}=i\widehat{\mathbb{R}}$ decomposes as $G_{0}g_{0}\cup\tau^{\prime}G_{0}g_{0}$. Since $G_{0}L\subset L$, the second point of the statement follows. ∎
Theorem B.
[…]
  1. i.
    $K_{n}$ is the stabilizer of $\Sigma$ in $\Gamma_{n}$.
  2. ii.
[…]
Equivalently, $K_{n}=K_{n}^{+}\cup K_{n}^{-}$ and $K_{n}^{\prime}=K_{n}^{+}\cup K_{n}^{-}\cup L_{n}^{+}\cup L_{n}^{-}$.
Proof.
Let $g$ be the matrix
$g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in\mathrm{PSL}_{2}(\mathbb{C}).$
Denote by $\sigma$ and $\sigma^{\prime}$ the automorphisms of $\mathrm{PSL}_{2}(\mathbb{C})$ given by conjugation by $\tau$ and $\tau^{\prime}$, respectively. Notice that $\sigma^{\prime}(g^{t})=g^{-1}$, so it follows that $g\cdot\Sigma=\Sigma$ if and only if $\sigma(g^{t})\cdot\widehat{\mathbb{R}}=\widehat{\mathbb{R}}$, and $g\cdot\mathbb{S}^{1}=i\widehat{\mathbb{R}}$ if and only if $\sigma(g^{t})\cdot\widehat{\mathbb{R}}=\mathbb{S}^{1}$. On the other hand, the quantities $\alpha$, $\beta$ and $\beta^{\prime}$ defined in (1) for $\sigma(g^{t})$ are the following:
$\alpha=i(a\overline{d}+c\overline{b}),\quad\beta=2\mathrm{Re}(b\overline{d}),\quad\beta^{\prime}=2\mathrm{Re}(a\overline{c}).$
It holds that $m(g)=0$ if and only if $\beta=\beta^{\prime}$. Also, this implies that $|\alpha|^{2}+|\beta|^{2}=1$. We consider now two cases:
  1. i.
    If $g\in\Gamma_{n}$, then $\beta,\beta^{\prime}\in 2\mathbb{Z}$, so we conclude that $\beta=\beta^{\prime}=0$ and $|\alpha|=1$. Therefore, $\sigma(g^{t})\cdot\widehat{\mathbb{R}}$ is a line that goes through 00 and has as orthogonal vector $\alpha$ (see Proposition 1.3). Notice also that $\alpha\in i\mathbb{Z}[\sqrt{-n}]$. If $n>1$, then $\alpha\in\{-i,i\}$, so we get $\sigma(g^{t})\cdot\widehat{\mathbb{R}}=\widehat{\mathbb{R}}$. If $n=1$, $\sigma(g^{t})\in\Gamma_{1}$. By Proposition 1.2 $,\sigma(g^{t})\cdot\widehat{\mathbb{R}}$ is tangent to $\widehat{\mathbb{R}}$, so they must be the same line.
  2. ii.
[…]
The rest of the statement follows from Lemma 2.1. ∎
Remark 2.2.
Whereas $K_{n}$ is always a subgroup of $\Gamma_{n}$, $K_{n}^{\prime}\subset\Gamma_{n}^{\prime}$ is not, since it is not closed under products neither taking inverses.
Remark 2.3.
The theorem does not apply for $K_{3}^{\prime}$. Notice that $\mathcal{O}_{-3}=\mathbb{Z}[\xi_{3}]$ where $\xi_{3}$ denotes a primitive $3$-root of the unit. A matrix as simple as:
$u=\begin{bmatrix}\xi_{3}&0\\ 0&\overline{\xi_{3}}\end{bmatrix}$
will be in $K_{3}^{\prime}$ but not in $K_{3}\cup L_{3}$. Also, since $m$ is right $\mathrm{PSU}(2)$-equivariant, if we pick any $g\in K_{3}$ then $gu\in K_{3}^{\prime}$, but this product will not be in $K_{3}\cup L_{3}$ in general.
3. Proof of the Cotlar Identity
The sets $K_{n}^{+}$ and $K_{n}^{-}$ defined in (4) verify certain relations related to the invariance of $m$: $\tau^{\prime}K_{n}^{+}=K_{n}^{+}\tau^{\prime}=K_{n}^{-}.$ These identities, together with the fact that $K_{n}^{+}$ is a subgroup of $\Gamma_{n}$, implies easily that:
$K_{n}^{+}K_{n}^{-},\,K_{n}^{-}K_{n}^{+}\subset K_{n}^{-}\quad\text{and}\quad K_{n}^{-}K_{n}^{-}\subset K_{n}^{+}.$
We claim now that, because of these inclusions, the function $\chi\colon K_{n}\to\mathbb{T}^{1}$ defined as:
$\chi(g)=\begin{cases}1&\text{if }g\in K_{n}^{+},\\ -1&\text{if }g\in K_{n}^{-},\end{cases}$
is a character. The following three lemmas prove that ${\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{\Gamma_{n}}}$ is left $(K_{n},\chi)$-equivariant.
Lemma 3.1.
Let $g\in\mathrm{PSL}_{2}(\mathbb{C})$ and let $r_{1}(g)$ and $r_{2}(g)$ denote the first and second rows of $g$, respectively. There exist an unitary matrix $u\in\mathrm{PSU}(2)$ such that:
$g=\begin{bmatrix}s^{-1}&s^{-1}t\\ 0&s\end{bmatrix}u,$
with $s=|r_{2}(g)|$ and $t=\langle r_{1}(g),r_{2}(g)\rangle$, where the bracket represents the scalar product in $\mathbb{C}^{2}$.
Proof.
This is just an explicit statement of the $\mathrm{ANK}$ decomposition for $\mathrm{PSL}_{2}(\mathbb{C})$. It can be proven directly as follows. Let $u$ be the (only) unitary matrix such that $r_{2}(g)u^{*}=(0,s)$ with $s>0$. Thus, $s=|r_{2}(g)|$. On the other hand, using that $\det(gu^{*})=1$, we get that $r_{1}(g)u^{*}=(s^{-1},\omega)$ for some $\omega\in\mathbb{C}$. This $\omega$ can be computed using that $\omega=s^{-1}\langle r_{1}(gu^{*}),r_{2}(gu^{*})\rangle=s^{-1}\langle r_{1}(g),r_{2}(g)\rangle$, which is the definition of $s^{-1}t$. ∎
Lemma 3.2.
For any $g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in\mathrm{PSL}_{2}(\mathbb{C})$, it holds that:
$\mathrm{Im}(b\overline{c}-a\overline{d})^{2}-4\mathrm{Re}(a\overline{c})\mathrm{Re}(b\overline{d})\leq 1.$
Moreover, if $g\in\Gamma_{n}$, then the right-hand side of the above inequality can be improved to 00.
Proof.
Same computations as in the proof of [4, Lemma 5.3] shows that the left-hand side of the above expression can be written as $p(X)=-4X(1+X)$, where $X=na_{2}d_{2}+b_{1}c_{1}$. This proves the statement for $g\in\mathrm{PSL}_{2}(\mathbb{C})$. If $g\in\Gamma_{n}$, then $X$ is an integer and therefore $p(X)\in 4\mathbb{Z}$, which proves the second part of the statement. ∎
Lemma 3.3.
The symbol ${\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{\Gamma_{n}}}$ is right $K_{n}$-invariant and left $(K_{n},\chi)$-equivariant.
Proof.
It is immediate to check that $m(\omega g)=-m(g).$ On the other hand, $K_{n}^{+}$ and $K_{n}^{-}$ are contained respectively in $G_{0}$ and $\omega G_{0}$, where $G_{0}$ is the group defined in the Introduction by (2). Since $m$ is invariant by the left action of $G_{0}$, it follows that ${\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{\Gamma_{n}}}$ is left $(K_{n},\chi)$-equivariant.
For the right invariance, let’s take $g\in\Gamma_{n}$ and $h\in K_{n}$. If $g\in K_{n}$, it is immediate that $m(gh)=m(g)=0$, so we rule out this case. Let’s write $g$ and $h$ as
$g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\quad\text{and}\quad h=\begin{bmatrix}s^{-1}&s^{-1}t\\ 0&s\end{bmatrix}u,$
[…]
$\begin{split}\mathrm{Re}\langle r_{1}(gh),r_{2}(gh)\rangle&=\mathrm{Re}(a\overline{c})(1+(\mathrm{Im}t)^{2})s^{-2}+\mathrm{Re}(b\overline{d})s^{2}+\mathrm{Im}(b\overline{c}-a\overline{d})\mathrm{Im}t\\ &=\begin{bmatrix}s^{-1}\mathrm{Im}t&s\end{bmatrix}\begin{bmatrix}2\mathrm{Re}(a\overline{c})&\mathrm{Im}(b\overline{c}-a\overline{d})\\ \mathrm{Im}(b\overline{c}-a\overline{d})&2\mathrm{Re}(b\overline{d})\end{bmatrix}\begin{bmatrix}s^{-1}\mathrm{Im}t\\ s\end{bmatrix}\\ &\qquad+s^{-2}\mathrm{Re}(a\overline{c})\end{split}$ (5)
The Lemma 3.2 says that the determinant of the matrix in (5) is always non-negative. Therefore, this matrix will be semidefinite positive if $\mathrm{Re}(a\overline{c})\geq 0$ and $\mathrm{Re}(b\overline{d})\geq 0$ and semidefinte negative otherwise. In both cases, it implies that $m(gh)m(g)\geq 0$. Since $gh\not\in K_{n}$, it follows that $m(gh)=m(g)$, proving the statement. ∎
In the proof of the Theorem A, we will make use of two inequalities that we introduce now as two independent lemmas.
Lemma 3.4.
Let $g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in\Gamma_{n}$. Then
$m(g)m(g^{t})\mathrm{Re}(a\overline{d}+b\overline{c})\geq 0,$
where $g^{t}$ denotes the transpose of $g$.
Proof.
If $g$ or $g^{t}$ are in $K_{n}$, the result is immediate. If they are not, we know $m(g)$ has the same sign as $\mathrm{Re}(a\overline{c})$ or $\mathrm{Re}(b\overline{d})$, depending on which one is non-zero. We’ll suppose that both $\mathrm{Re}(a\overline{c})$ and $\mathrm{Re}(a\overline{b})$ are non-zero, since the rest of the cases comes from applying this one to $\tau^{\prime}g$, $g\tau^{\prime}$ or $\tau^{\prime}g\tau^{\prime}$.
Under this hypothesis, the statement is equivalent to
$\mathrm{Re}(a\overline{c})\mathrm{Re}(a\overline{d})\mathrm{Re}(a\overline{d}+b\overline{c})\geq 0.$
From the proof of [4, Proposition 5.8] we know that the left-hand side of the inequality equals $p(X)=(AX+B)(2X+1)$ with $A=n(a_{1}^{2}+a_{2}^{2})$, $B=na_{2}^{2}$ and $X=b_{1}c_{1}+na_{2}d_{2}$. Since $X$ is an integer and the roots of the polynomial $p$ have modulus lesser or equal than $1$, we conclude the statement. ∎
Lemma 3.5.
For any $g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in\mathrm{PSL}_{2}(\mathbb{C})$, it holds that $\mathrm{Re}(a\overline{c})\mathrm{Re}(b\overline{d})\geq-\frac{1}{4}$. Moreover, if $g\in\Gamma_{n}$, the right-hand side of the inequality can be improved to 00.
Proof.
Suppose that $\mathrm{Re}(a\overline{c})\mathrm{Re}(b\overline{d})<0$. Then, by multiplying the equation $ad-bc=1$ by $\overline{cd}$ and taking real part, we get that:
$|\mathrm{Re}(b\overline{d})||c|^{2}+|\mathrm{Re}(a\overline{c})||d|^{2}=|\mathrm{Re}(c\overline{d})|\leq|c||d|.$
Now we claim that any positive numbers $x,y,\alpha,\beta>0$ satisfying
$\alpha x^{2}+\beta y^{2}\leq xy$ (6)
must verify $\alpha\beta\leq\frac{1}{4}$. To prove the claim, just notice that (6) is equivalent to $\alpha u^{2}-u+\beta\leq 0$ with $u=x/y$, and this can only happen if the discriminant $1-4\alpha\beta$ is greater than or equal to 00.
If $g\in\Gamma_{n}$, then both $\mathrm{Re}(a\overline{c})$ and $\mathrm{Re}(b\overline{d})$ are integers, so $\mathrm{Re}(a\overline{c})\mathrm{Re}(b\overline{d})$ must be indeed non-negative. ∎
Proof of Theorem A.
We are going to prove that the symbol ${\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{\Gamma_{n}}}$ satisfies (Cotlar) relative to $K_{n}$, that is:
$(m(g^{-1})-m(h))(m(gh)-m(g))=0,\text{ for all }g\in\Gamma_{n}\setminus K_{n}\text{ and }h\in\Gamma_{n}.$
If $h\in K_{n}$, the equality follows from the right $K_{n}$-invariance of $m$ proven in Lemma 3.3. Now, suppose that $h\not\in K_{n}$ and $m(g^{-1})\neq m(h)$. We have to prove that $m(gh)=m(g)$. Since the hypothesis $m(g^{-1})\neq m(h)$ implies that $gh\not\in K_{n}$, it suffies to prove that $m(gh)m(g)\geq 0$. We write:
$g=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\quad\text{and}\quad h=\begin{bmatrix}s^{-1}&s^{-1}t\\ 0&s\end{bmatrix}u,$
using Lemma 3.1 to decompose $h$ into an upper-triangular matrix and an unitary one. Now, a computation shows that:
$\begin{split}m(gh)m(g)&=\operatorname{sign}\Big{(}\mathrm{Re}(a\overline{c}+b\overline{d})\mathrm{Re}(a\overline{c})s^{-2}(1+(\mathrm{Re}t)^{2})\\ &\qquad+\mathrm{Re}(a\overline{c}+b\overline{d})\mathrm{Re}(a\overline{d}+b\overline{c})\mathrm{Re}t\\ &\qquad+\mathrm{Re}(a\overline{c}+b\overline{d})\big{[}\mathrm{Re}(a\overline{c})s^{-2}(\mathrm{Im}t)^{2}+\mathrm{Re}(b\overline{d})s^{2}+\mathrm{Im}(b\overline{c}-a\overline{d})\mathrm{Im}t\big{]}\Big{)}\\ &=\operatorname{sign}\Big{(}(\mathrm{I})+(\mathrm{II})+(\mathrm{III})\Big{)}\end{split}$
Now notice that $(\mathrm{I})$ is non-negative because of Lemma 3.5 and the fact that $s>0$. Also, $(\mathrm{II})$ is non-negative because $\mathrm{Re}t$ has the same sign as $m(h)=-m(g^{-1})=m(g^{t})$, so we can apply Lemma 3.4. Finally, $(\mathrm{III})$ is non-negative because of Lemma 3.2, which implies that each factor of the product has the same sign as $m(g)$ or is zero. ∎
Remark 3.6.
We still don’t know if the Fourier multiplier given by ${\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{\Gamma_{n}^{\prime}}}$ is bounded or not in $L_{p}(\mathcal{L}\,\Gamma_{n}^{\prime})$, but what can be proven is that this symbol do not verify a Cotlar identity as in Theorem 1.1 with respect to any possible subgroup of $\Gamma_{n}^{\prime}$. To see this, suppose that ${\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{\Gamma_{n}^{\prime}}}$ is $(G_{0},\chi)$-equivariant for some subgroup $G_{0}\leq\Gamma_{n}^{\prime}$ and some character $\chi$ on $G_{0}$. We claim that $G_{0}\cap L_{n}^{+}=\varnothing$. Firstly, notice that for any $l\in L_{n}^{+}$ and $g\in\mathrm{PSL}_{2}(\mathbb{C})$, it holds that:
$m(lg)=\operatorname{sign}\left(|r_{1}(g)|-|r_{2}(g)|\right)$
where $r_{1}(g)$ and $r_{2}(g)$ denotes the first and second rows of $g$ as complex vectors in $\mathbb{C}^{2}$. Let $a\colon\Gamma_{n}^{\prime}\to\Gamma_{n}^{\prime}$ be the map that permutes the two rows of a matrix and multiplies the first column by $-1$. If $G_{0}\cap L_{n}^{+}\neq\varnothing$, by the formula above it would hold that for any $h\in\Gamma_{n}^{\prime}$ and any $l\in G_{0}\cap L_{n}^{+}$:
$m(lh)=\chi(l)m(h)=\chi(l)m(a(h))=m(la(h))=-m(lh),$
which is of course impossible. On the other hand, fix an $l\in L^{+}_{n}$ whose inverse is not in $K_{n}^{\prime}$. Then in order to Cotlar identity to hold, one needs that:
$m(l^{-1})=m(h),\quad\text{for any }h\in\Gamma_{n}^{\prime}\text{ such that }m(lh)\neq 0.$
Pick $h\in\Gamma_{n}^{\prime}$ any element which verifies this equation. Let $h^{\prime}$ be given by $h^{\prime}=\tau^{\prime}h\tau^{\prime}$, where $\tau^{\prime}$ is the matrix defined at the beginning of Section 2. Notice that $m(lh^{\prime})=-m(lh)\neq 0$, but $m(h^{\prime})=-m(h)$. Therefore Cotlar identity must fail when applied to $l$ and $h^{\prime}$.
4. Addenda: another $L_{p}$-bounded multiplier on Bianchi groups
Our initial choice of hyperbolic plane $\Sigma=\{\omega\in\mathbb{H}^{3}\colon\mathrm{Re}(\omega)=0\}$ was motivated by the authors of [4] proving that the corresponding multiplier $m$ satisfies the Cotlar identity in $\mathrm{SL}_{2}(\mathbb{Z})$ and $\mathrm{SL}_{2}(\mathbb{Z}[i])$. However, we have seen that such an identity fails for $m$ when restricted to a general Bianchi group $\Gamma_{n}^{\prime}$. This failure is connected to the geometry of circles in the orbit of $\partial\Sigma$ under the action of $\Gamma_{n}^{\prime}$, so it is natural to ask for multipliers induced by planes with a better-behaved boundary. Concretely, we are going to consider now the multiplier $\widetilde{m}$ induced by the hyperplane $\widetilde{\Sigma}=\{\omega\in\mathbb{H}^{3}\colon\mathrm{Im}(\omega)=0\}$, and the partition $\widetilde{\Sigma}^{+}=\{\omega\in\mathbb{H}^{3}\colon\mathrm{Im}(\omega)>0\}$, $\widetilde{\Sigma}^{-}=\{\omega\in\mathbb{H}^{3}\colon\mathrm{Im}(\omega)<0\}$.
Indeed, we claim that $m$ will satisfy the Cotlar identity on $\mathrm{SL}_{2}(\mathbb{Z}[i])$ if and only if $\widetilde{m}$ does so. Let $\sigma$ be the automorphism of $\mathrm{PSL}_{2}(\mathbb{C})$ given by conjugation by $\tau$, where $\tau$ is the matrix defined at the beginning of Section 2. A simple computation shows that:
$\widetilde{m}(\sigma(g))$ $=\mathrm{sign}\,\mathrm{Im}\left[(aj+bi)\overline{(-cij+d)}\right]$
$=\mathrm{sign}\,\mathrm{Re}\left(a\overline{c}+b\overline{d}\right)=m(g).$
The automorphism $\sigma$ leaves $\mathrm{SL}_{2}(\mathbb{Z}[i])$ invariant, so $\sigma$ restricts to an automorphism of this group, proving our claim. Therefore, this point of view also generalizes the results of [4] in a different way than we did before.
Lemma 4.1.
For any $n\geq 1$ with $n\neq 3$, the set $\widetilde{K}_{n}=\{g\in\Gamma_{n}^{\prime}\colon\widetilde{m}(g)=0\}$ coincides with the stabilizer of $\widetilde{\Sigma}$ under the action of $\Gamma_{n}^{\prime}$. Indeed, $\widetilde{m}(g)$ can be written as:
$\widetilde{m}(g)=\begin{cases}1&\text{if }g\cdot\widehat{\mathbb{R}}\text{ lies on the upper half-plane},\\ -1&\text{if }g\cdot\widehat{\mathbb{R}}\text{ lies on the lower half-plane},\\ 0&\text{otherwise}.\end{cases}$
Proof.
Notice first that two hyperbolic planes $\Sigma_{1}$ and $\Sigma_{2}$ verify $\Sigma_{1}\cap\Sigma_{2}\neq\emptyset$ if and only if $\partial\Sigma_{1}$ and $\partial\Sigma_{2}$ intersect in at least two points. If $g\in\Gamma_{n}^{\prime}$ and $\widetilde{m}(g)=0$, it is because $g\cdot\widetilde{\Sigma}\cap\widetilde{\Sigma}\neq\emptyset$. Since $\partial\widetilde{\Sigma}=\widehat{\mathbb{R}}$, this implies that $g\cdot\widehat{\mathbb{R}}$ and $\widehat{\mathbb{R}}$ intersect in at least two points, but because of Proposition 1.2 it means that they are the same generalized circle. Therefore $g\cdot\widehat{\mathbb{R}}=\widehat{\mathbb{R}}$ and $g\cdot\widetilde{\Sigma}=\widetilde{\Sigma}$. The rest of the statement follows similarly. ∎
Lemma 4.2.
The multiplier $\widetilde{m}$ restricted to $\Gamma_{n}^{\prime}$ is right $\widetilde{K}_{n}$-invariant and left $(\widetilde{K}_{n},\chi)$-equivariant, for some character $\chi$ on $\widetilde{K}_{n}$.
Proof.
Both the invariance and the equivariance follow easily from Lemma 4.1. just taking the character $\chi$ defined as:
$\chi(g)=\begin{cases}1&\text{if }g\text{ send the upper half-plane to itself},\\ -1&\text{otherwise}.\end{cases}$
Theorem 4.3.
The symbol $\widetilde{m}$ satisfies the Cotlar identity on every Bianchi group $\Gamma_{n}^{\prime}$ with $n\neq 3$.
Proof.
Because of Lemma 4.2, it is enough to prove that
$(\widetilde{m}(g^{-1})-\widetilde{m}(h))(\widetilde{m}(gh)-\widetilde{m}(g))=0,\text{ for all }g,h\in\Gamma_{n}^{\prime}\smallsetminus\widetilde{K}_{n}.$
Let’s suppose that $\widetilde{m}(g^{-1})\neq\widetilde{m}(h)$. Then, without lost of generality, we can suppose that there is a generalized circle $C_{g}$ in the upper half-plane and a generalized circle $C_{h}$ in the lower half-plane such that $g\cdot C_{g}=\widehat{\mathbb{R}}$ and $h\cdot\widehat{\mathbb{R}}=C_{h}$. Since $\widehat{\mathbb{R}}$ and $C_{h}$ lie both on the exterior of $C_{g}$, they are mapped into the same half-plane by $g$. That is, $g\cdot C_{h}=gh\cdot\widehat{\mathbb{R}}$ and $g\cdot\widehat{\mathbb{R}}$ lie on the same side of $\widehat{\mathbb{C}}\smallsetminus\widehat{\mathbb{R}}$, and therefore $\widetilde{m}(gh)=\widetilde{m}(g)$. ∎
Acknowledgments
The author was partially supported by pre-doctoral scholarship PRE2020-093245, the Severo Ochoa Grant CEX2023-001347-5(MICIU), PIE2023-50E106(CSIC), and PID2022-141354NB-I00(MICINN). He would also like to thank his advisor, Adrián González-Pérez, for pointing him out the problem, as well as for his support and insightful conversations.